OpenStudy (anonymous):

a cylindrical tree trunk is 14 yards high from the ground up to the lowest branch, and it measures 4 yards around. what is the volume of wood in this section of the trunk? A. 224pi yd^3 B. 56/pi yd^3 C. 56pi yd^3 D. 28 yd^3 E. 28pi yd^3

OpenStudy (campbell_st):

use the circumference information to find the radius... \[4 = 2\pi r\] solve for r then the volume is \[V = \pi r^2 h\] you know h = 14 and found r earlier... just substitute to find the volume of wood

OpenStudy (anonymous):

thx men